Circular No. 8065

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ad Further to IAUC 8062, T. Puckett and J. Newton report the discovery

of an apparent supernova (mag 17.8) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Feb. 3.47 UT (and confirmed by Newton on a frame taken on Feb. 4.32 with a 0.40-m reflector). The new object is located at α = 14^{h}56^{m}28^{s}.07, δ = +45^{o}21′09′′.2 (equinox 2000.0), which is 7′′.4 east and 4′′.2 south of the center of its apparent host galaxy. SN 2003ad is not present on images taken by Puckett on 2001 July 14 and 2002 Feb. 14 (limiting mag ∼ 20.0) or on Palomar Sky Survey plates taken on 1994 June 6, Apr. 16 (limiting mag ∼ 21.0), and on 1955 Mar. 25.

SUPERNOVAE 2003X, 2003ab, AND 2003ac R. J. Foley, B. Swift, J. Graham, D. S. Wong, R. Chornock, and A. V.

Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm), obtained on Feb. 4 UT with the Shane 3-m telescope at Lick Observatory, reveals that SN 2003X (IAUC 8062) is of type Ia, resembling SN 1994D roughly 1 month past maximum brightness. SN 2003ab (IAUC 8064) is of type II, within a few weeks after the explosion. SN 2003ac (IAUC 8064) is also of type II, but the rather noisy spectrum resembles most closely that of the type-IIb supernova 1987K (Filippenko 1988, A.J. 96, 1941); further spectral monitoring may reveal emerging He I lines.

COMET C/2003 B1 (SOHO) Further to IAUC 8032, D. Hammer reports measurements for a Meyer-

group comet found on SOHO website C2 and C3 images by X.-M. Zhou (reduced astrometry and orbital elements on MPEC 2003-C23).

2003 UT  α_{2000}  δ_{2000}
Jan. 16.814  19^{h}59.^{m}5  −20^{o}22′

COMET C/2002 Y1 (JUELS-HOLVORCEM) Visual m_{1} estimates: Jan. 8.83 UT, 11.8 (Y. Nagai, Yamanashi, Japan,

0.32-m reflector); 13.78, 12.2 (O. Miyazaki, Ishioka, Ibaraki, Japan, 0.32-m reflector); 26.07, 10.3 (J. Carvajal, Toledo, Spain, 0.45-m reflector); 27.25, 9.9 (B. H. Granslo, Fjellhamar, Norway, 0.15-m reflector); 29.03, 8.5 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars).

2003 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green